an object O is stuck on the surface of a transparent solid sphere of radius 20cm. The refractive index of the sphere such that rays from the object after refraction from the opposite side emerge as a parallel beam is?

Grade:11

1 Answers

Srihari
6 years ago
Mu2/infinity -Mu1/-40=Mu2-Mu1/-20 ......... Mu1/40=1-M1/-20 after solving the above equation we get Mu=2